OpenStudy (anonymous):

solve the equation (7s+4)(3s-5)=0

OpenStudy (tkhunny):

The premise is this: 1) We are multiplying two things. 2) The product of the two things is zero (0). 3) What do we know about at least one of the two things? One MUST be zero. 4) Find where the things are zero.

OpenStudy (anonymous):

so 4/7, 5/3?

OpenStudy (imtant):

21s^2-35s+12s-20=0 21s^2-23s-20=0 Then factorise it,there are two values of s..

OpenStudy (tkhunny):

s = -4/7 or s = 5/3 That's all there is to it!

OpenStudy (anonymous):

thats right its negative! thanks!

OpenStudy (imtant):

Ya,it is..